Enter your answer in the provided box. sample of chlorine gas is confined in a 8.01-L container at 626 torr and 59.0°C. How many moles of gas are in the sample? mol Cl2
Answer #1

PV = nRT

where, P = 626 torr = 0.824 mole

V = 8.01 L

n = number of moles

R = Gas constant

T = 59.0 + 273 = 332 K

0.824 * 8.01 = n * 0.0821 * 332

6.60 = n * 27.3

n = 6.60 / 27.3 = 0.242 mole

Therefore, the number of moles of gas = 0.242